OpenStudy (anonymous):

Calculate the pull on the root of a turbine blade due to centrifugal force when the rotor is rotating at 300 rpm, the mass of the blade is .2kg and the radius from the axis of the rotor to the center of gravity of the block is 460mm.

OpenStudy (anonymous):

@eliassaab @Kainui

OpenStudy (anonymous):

@iambatman sir lets continue please

OpenStudy (shamim):

Centrifugal force=mv^2/r=m*omega^2*r=? Given m=.2kg r=460mm=.46m Omega=2*pi*n n=300rpm=300/60hz

OpenStudy (shamim):

Feel free to ask for more clarification